OpenStudy (anonymous):

Using Boyle's Law (assuming constant pressure) A sample of oxygen gas occupies a volume of 250 mL at 740 torr pressure. What volume will it occupy at 800 torr pressure?

OpenStudy (anonymous):

I only know the full ideal gas law: PV=nrT n = PV/rT We know that the number of molecules will stay the same so: PV/rT = PV/rT (in the two situations) PV = PV (gas constant and temp are the same) 740 * 250 = 800V V = 185000/800 V = 231.25mL Never heard of the torr before. Interesting measurement

OpenStudy (anonymous):

... uh 760mmHg = 760 torr=1atm

OpenStudy (joannablackwelder):

Boyle's law is P1V1=P2V2 so 740*250=800*V2. Solve for V2.